# Unified Gaussian Primitives for Scene Representation and Rendering

### v1

@@ 1. Introduction
The adaptability of Gaussians to complex geometry, as demonstrated by 3DGS, inspires us to develop a Gaussian-based rendering primitive not just for radiance fields, but for general-purpose light transport and appearance modeling. Crucially, we design each Gaussian to be an atomic primitive, meaning that we do not explicitly simulate volumetric random walks within it. This concept is analogous to a triangle, the primitive for mesh-based representations. This approach contrasts with defining the spatially varying extinction coefficient as a sum of Gaussians, where sub-Gaussian scattering still requires explicit tracking. To achieve this, we combine the Gaussians with a non-exponential, linear transmittance model (Vicini et al., 2021), such that the free-flight distributions of Gaussians can be disentangled. The linear transmittance model allows our primitive to adapt to both hard surfaces and aggregated elements that exhibit semi-transparency. Each Gaussian primitive can thus be assigned with its own phase function, enabling meaningful appearance modeling and authoring. Using our primitives, a scene becomes a novel kind of heterogeneous, non-exponential volume. We derive efficient Monte Carlo operations, such as free-flight distribution sampling, for the volume to be rendered by Monte Carlo path tracing with full global illumination.
@@ 2. Related Work
Recently, Kerbl et al. (2023) develop 3D Gaussian splatting (3DGS) that extends the EWA volume splatting framework (Zwicker et al., 2001a) to be differentiable and uses 3D Gaussians to optimize and render radiance fields. 3DGS achieves state-of-the-art reconstruction quality and offers significantly faster rendering speed compared to previous NeRF approaches. Since its debut, 3DGS has inspired a number of Gaussian-based representations with different focuses, such as for mesh reconstruction (Huang et al., 2024; Guédon and Lepetit, 2023), avatar rendering (Saito et al., 2023), and inverse rendering (Gao et al., 2023). While not modeling the full light transport, 3DGS demonstrates the effectiveness of anisotropic Gaussians in adapting to complex geometries, especially thin structures.
@@ 3.2. Volumetric Light Transport
where $\sigma_{\mathrm{t}}$ is the spatially varying extinction coefficient, which intuitively controls the density of the volume. Note that $\mathcal{T}$ and $\mathcal{P}$ only differ by $\sigma_{\mathrm{t}}$. This is coincidentally due to the unique property of the exponential function being invariant under differentiation. Because $\mathcal{P}$ includes the nonlinear exponential function, the exponential free-flight PDF is not additive. This introduces a theoretical challenge if one attempts to define a volume as the sum of “exponential Gaussian primitives”. To see this, suppose we define the extinction coefficient of an exponential volume as $\sigma_{\mathrm{t}}(x)=\sum_{k}G_{k}(x)$. The total free-flight PDF is then different from the sum of free-flight PDFs of individual Gaussians:
(6) $\begin{split}&\Big{(}\sum_{k}G_{k}(x_{t})\Big{)}\exp\Big{(}-\int_{0}^{t}\Big{(}\sum_{k}G_{k}(x_{t^{\prime}})\Big{)}\mathrm{d}{t^{\prime}}\Big{)}\neq\\ &\sum_{k}G_{k}(x_{t})\exp\Big{(}-\int_{0}^{t}G_{k}(x_{t^{\prime}})\mathrm{d}{t^{\prime}}\Big{)}.\end{split}$
This implies that it is impossible to define the property of each Gaussian in isolation and “assemble” them together to form the final volume. Instead, the entire continuous volume must be treated in its entirety. Therefore, it is not suitable to define a Gaussian as a rendering primitive within the exponential formulation.
@@ 4. Linear Transmittance Gaussian Primitives
Our goal is to define a general-purpose volumetric rendering primitive based on 3D Gaussian distribution. We begin by analyzing the feasibility and requirements for defining such a primitive. 3DGS has convincingly demonstrated the advantages of anisotropic 3D Gaussians for adapting to complex shapes. However, to be truly usable in light transport, we need to define how these Gaussians interact with light. This includes the attenuation of light, controlled by the free-flight distribution or transmittance, and the scattering (or absorption) of light, controlled by the appearance or phase function. For the primitives to be practically valuable in modeling and rendering applications, the following properties are desirable:
It should support hard surfaces. Just like radiance fields, this capability allows the primitive to model a much wider range of objects, instead of being limited to traditional “volume-like” objects such as clouds and smoke. Unlike radiance fields, however, exponential volumes with full light transport simulation are prone to artifacts such as leaking. To address this issue, we draw inspiration from the hybrid transmittance function by Vicini et al. (2021), as its linear component models negative correlation for hard surfaces.
Appearance should be defined on a per-primitive level. We treat primitives as the atomic elements of the scene, using appearance to abstracts away all sub-primitive scattering interactions. This decision benefits efficiency by avoiding explicit simulation of random walks inside each primitive. Additionally, since fitting any non-trivial geometry often results in many overlapping Gaussians, the appearance of overlapped regions should be well-defined for meaningful authoring and editing. As previously discussed, this is not achievable with exponential transport (Eq. 6).
@@
where the saturating distance $t_{\mathrm{sat}}$ is the ray travel distance such that $\mathcal{T}(x,x_{t_{\mathrm{sat}}})=0$. Fig. 2 illustrates the above definitions in 2D flatland. It is clear from Eq. 10 that the total free-flight PDF is simply the sum of the PDFs of individual primitives. This implies that when integrating § 3.2, we can sample an individual primitive and compute its contribution separately from other primitives. In other words, the appearance of a primitive is decoupled from other overlapping primitives. The linear transmittance improves the ability to model negatively correlated hard surfaces. On the other hand, it does not compromise the ability to model objects with more stochastic characteristics such as foliage. Indeed, as we will demonstrate in § 7.2, with suitable optimization, the final heterogeneous transmittance can approach an arbitrary target.
@@ 4.1. Primitive Operations
3D Gaussians support several key operations for volumetric rendering that are either in closed forms or can be computed efficiently. These operations serve as the building blocks of our scene rendering algorithms, introduced in § 5. We further define the corresponding appearance for a Gaussian primitive in § 6. Notably, our Gaussian primitive exhibits close resemblance to how a triangle serves as the primitives in mesh-based representations.
@@ 5.1. Free-flight Distribution Sampling
Refer to caption Refer to caption Refer to caption
@@ 6. Appearance
The linear transmittance model ensures that the probability of a ray scattered by each Gaussian primitive is simply additive. It allows us to conveniently define the appearance of each primitive individually. In this work, we define appearance models at the primitive level, conceptually abstracting away all sub-primitive scattering interactions. This design decision eliminates the need to explicit simulate random walks inside each primitive, thereby making rendering more efficient.
A Gaussian primitive can simply be a part of a flat surface, or represents a collection of small surface elements with different orientations. Therefore, a useful appearance model for the primitive should incorporate this aggregated effect while still being compatible with simple flat scenarios. Inspired by the microflake theory (Jakob et al., 2010; Heitz et al., 2015), we describe the orientations by a normal distribution function (NDF) $D(n)$ and its visible normal distribution function (VNDF) $D_{\omega_{o}}(n)$ when conditioned by a viewing direction $\omega_{o}$:
@@
The base surface BSDF can be arbitrary valid BSDF with standard evaluation and importance sampling procedures. In this work, we primarily work with the Disney principled BSDF (Burley, 2015) due to its capability to model a wide range of materials. The Disney principled BSDF mainly consists of a microfacet specular component and a diffuse component with empirical retroflection. We refer readers to Burley (2015) for the full definition.
@@
We follow the original importance sampling strategy by (Heitz et al., 2015). Given a view direction $\omega_{o}$, we first sample the VNDF to generate a sample $n$. We then sample an incident direction $\omega_{i}$ using the existing strategy of the base BSDF. Since the VNDF sampling for SGGX is perfect, the sample weight is simply that of the base BSDF sampling.
@@ 7.2. Transmittance Optimization
Transmittance Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Tr. Loss Refer to caption Refer to caption
@@ 8. Applications and Results
In Fig. 1, we demonstrate the versatility of our Gaussian primitive to represent objects with a wide range of geometric and material characteristics. The Dressing Table scene is modeled entirely by our primitives and contains parts that are acquired in different ways. The room, table, dragon, and logo are converted from meshes, while the plants, candle set, and blanket are converted from 3DGS reconstruction. Additionally, several objects, such as the mirror and neon light bars, are modeled analytically. Our Gaussian primitives can adapt to different shapes, including flat surfaces and thin fibers, thanks to their anisotropic definition. The volumetric formulation naturally handles semi-transparency for dense, stochastic details.
@@
Crucially, our framework supports full unbiased global illumination with Monte Carlo (volumetric) path tracing. This is in contrast to rasterization approaches including 3DGS, and volumetric ray marching approaches such as NeRF. Both families of approaches are fundamentally limited to the first few dimensions of the path space, and are simply incapable of solving the infinite dimensional light transport integral. Fig. 1 includes two renders with drastically different lighting setups. In the “daylight” setup, the scene is lit by an area light and an environment light; in the “night” setup, the scene is lit by a logo, neon light bars, and candles, all modeled as emissive Gaussian primitives. The number of emissive primitives exceeds 10K combined. Both renders showcase various global illumination effects including soft shadows, color bleeding, and inter-reflection. Additional sequences with zoomed-in camera animations are provided in the supplemental video.
@@
In Fig. 11, we compare our representation to the traditional volume representation consisting of regular voxels. We perform a simple voxelization of our models by resampling the free-flight PDF. For each voxel, we evaluate all overlapping Gaussians at the voxel center. The phase function parameters are similarly resampled. The voxel grid is then linearly interpolated and rendered by ray marching. We evaluate the reconstruction quality using different voxel resolutions. Even using 8$\times$ more voxels than the number of Gaussians, the reconstruction quality is still significantly more inferior. This is expected because unlike Gaussian mixtures that can approximate signals at arbitrary frequency, regular grid sampling is limited by the Nyquist-Shannon sampling theorem and the resolution must be at least twice the signal bandwidth to avoid aliasing. It would require an impractical amount of storage to properly represent the thin structures common in vegetation. Conversely, Gaussian primitives are much more effective at capturing fine geometry details.
@@
Figure 11. Comparison between Gaussian models and voxelized models in different resolutions. Gaussian primitives are much more effective at representing sharp, detailed structure than voxels. Note that our current simple voxelization causes self-shadowing that leads to slight darkening. Direct illumination only.
@@
Garden Stump
[…]
(a) Original
[…]
(b) Lighting 1
[…]
(c) Lighting 2

### v2

@@ 1. Introduction
The adaptability of Gaussians to complex geometry, as demonstrated by 3DGS, inspires us to develop a Gaussian-based rendering primitive not just for radiance fields, but for general-purpose light transport and appearance modeling.
Crucially, we design our Gaussian primitive to be atomic, an indivisible entity, and we do not explicitly simulate sub-primitive random walks. A parallel can be drawn between our Gaussian primitive and a triangle as they are both the basic building block for the respective scene representations. Each Gaussian primitive can be assigned with its own per-primitive phase function, enabling meaningful appearance modeling and authoring. This approach contrasts with defining an exponential volume with spatially varying extinction coefficient as a sum of Gaussians, where sub-Gaussian multiple scattering is required to resolve appearance. We further combine the Gaussians with a non-exponential, linear transmittance model (Vicini et al., 2021) to better adapt to both hard surfaces and fuzzy, aggregated elements.
Using our primitives, a scene becomes a novel kind of heterogeneous, non-exponential volume. We derive efficient Monte Carlo operations, such as free-flight distribution sampling, for the volume to be rendered by Monte Carlo path tracing with full global illumination.
@@ 2. Related Work
Recently, Kerbl et al. (2023) develop 3D Gaussian splatting (3DGS) that extends the EWA volume splatting framework (Zwicker et al., 2001a) to be differentiable and uses 3D Gaussians to optimize and render radiance fields. 3DGS achieves state-of-the-art reconstruction quality and offers significantly faster rendering speed compared to previous NeRF approaches. Since its debut, 3DGS has inspired a number of Gaussian-based representations with different focuses, such as for level-of-detail rendering (Kerbl et al., 2024), mesh reconstruction (Huang et al., 2024; Guédon and Lepetit, 2023), avatar rendering (Saito et al., 2023), inverse rendering (Gao et al., 2023), and high-dimensional function fitting (Diolatzis et al., 2024). While not modeling the full light transport, 3DGS demonstrates the effectiveness of anisotropic Gaussians in adapting to complex geometries, especially thin structures. 3DGS relies on an approximate global depth sort operation that introduces inaccuracies such as popping. Recent works explore more accurate rendering methods than splatting, such as hierarchical rasterization (Radl et al., 2024) or ray tracing (Moenne-Loccoz et al., 2024). However, they are either limited to primary intersection or Whitted-style ray tracing. Our method supports Monte Carlo path tracing with arbitrary bounces and non-specular appearance.
@@ 3.2. Volumetric Light Transport
where $\sigma_{\mathrm{t}}$ is the spatially varying extinction coefficient, which intuitively controls the density of the volume. Note that $\mathcal{T}$ and $\mathcal{P}$ only differ by $\sigma_{\mathrm{t}}$. This is coincidentally due to the unique property of the exponential function being invariant under differentiation.
@@ 4. Linear Transmittance Gaussian Primitives
Our goal is to define a general-purpose volumetric rendering primitive based on 3D Gaussian distribution. We begin by analyzing the feasibility and requirements for defining such a primitive. 3DGS has convincingly demonstrated the advantages of anisotropic 3D Gaussians for adapting to complex shapes. However, to be truly usable in light transport, we need to define how these Gaussians interact with light. This includes the attenuation of light, controlled by the free-flight distribution or transmittance, and the scattering (or absorption) of light, controlled by the appearance or phase function. For the primitive to be practically valuable in modeling and rendering applications, the following properties are desirable:
Prioritization of Opaque Surfaces and Elements. Instead of true participating media such as clouds and smoke, we intend to use the primitive to represent the wider range of content consisting of opaque geometries. Objects may exhibit a spectrum of geometric characteristics from continuous surfaces to dense elements (which in aggregation produce a fuzzy look, but are nonetheless opaque individually). Exponential transport is not suitable in most of these cases, as the geometries are usually not independently distributed. Only when the elements are sufficiently decorrelated, the free-flight statistics may approach exponential in the far-field limit. However, this is also rarely the case: For a scene to reach high fidelity with sharp details, typically a single primitive only represents a surface patch or a small cluster of oriented elements. To address this issue, we draw inspiration from previous literature (Bitterli et al., 2018; Vicini et al., 2021) and incorporate the linear transmittance model to handle the negative correlation for opaque surfaces. We validate this design choice in Fig. 12.
Intuitive Appearance Definition. It is important to define the appearance of the primitive in a way that accommodates intuitive authoring and editing. Traditionally, the appearance of a volume is specified by the (single-scatteirng) phase function at each point in space, which governs how light is scattered at each scattering event. However, light transport in the volume undergoes multiple scattering, which contributes significantly to the final rendered image. This typically results in more saturated color and a softer look [^2] The relationship between the phase function and the apparent color of the volume is non-obvious and sometimes even ambiguous (Zhao et al., 2014; Wrenninge et al., 2017).. Multiple scattering is also costly as it requires explicit simulation of random walks inside each primitive. We argue that such per-point appearance definition is unnecessarily complex. In contrast, we propose to define appearance at a per-primitive level: the phase function describes the aggregated behavior for an entire primitive and abstracts away all sub-primitive scattering interactions. This also necessitates the need for a new kind of discrete free-flight sampling where a primitive is sampled each time, in contrast to the common continuous free-flight sampling where a collision point is sampled each time. Once a primitive is selected, light is scattered once according to the per-primitive phase function and then exits. Further self-intersections are ignored during this event. Conceptually, the role of our Gaussian primitive resembles that of a triangle in mesh-based representations: a basic, atomic entity for scene authoring with probabilistic scattering replacing deterministic intersection tests. Multiple scattering between primitives should still be tracked for global illumination. The conceptual differences between conventional volume path tracing and our formulation are illustrated in Fig. 2.
Figure 2. In conventional volume path tracing, a collision point is sampled at each time and the apparent appearance of a primitive depends on multiple scattering (left). Our method samples the entire primitive and uses the per-primitive phase function to scatters light once, essentially treating it as an indivisible entity.
@@
where the saturating distance $t_{\mathrm{sat}}$ is the ray travel distance such that $\mathcal{T}(x,x_{t_{\mathrm{sat}}})=0$. Fig. 3 illustrates the above definitions in 2D flatland. It is clear from Eq. 9 that the free-flight PDF is additive: the total free-flight PDF is simply the sum of the PDFs of individual primitives. This implies that when integrating §3.2, we can sample an individual primitive and compute its contribution separately from other primitives. In other words, the appearance of a primitive is decoupled from other overlapping primitives.
The linear transmittance improves the ability to model negatively correlated opaque geometries. On the other hand, it does not compromise the ability to model objects with more stochastic features such as foliage. Indeed, as we will demonstrate in §7.2, with suitable optimization, our final heterogeneous transmittance can fit these objects well. In Fig. 12, we will also validate that the linear transmittance model is overall superior to the exponential model for different variants of opaque geometries.
@@ 4.1. Primitive Operations
3D Gaussians support several key operations for volumetric rendering that are either in closed forms or can be computed efficiently. These operations serve as the building blocks of our scene rendering algorithms, introduced in §5. We further define the corresponding appearance for a Gaussian primitive in §6.
@@ 5.1. Free-flight Distribution Sampling
[…]
@@ 6. Appearance
The free-flight sampling in §5.1 allows us to probabilistically select an Gaussian primitive to scatter the light. The next ingredient is the primitive appearance that governs how light is scattered into different directions. As motivated in §4, we define appearance models at the primitive level and abstract away all sub-primitive scattering interactions.
A Gaussian primitive can represent a simple surface patch or a collection of small surface elements with different orientations. Therefore, a useful appearance model for the primitive should incorporate this aggregated effect while still being compatible with simple flat scenarios. Inspired by the microflake theory (Jakob et al., 2010; Heitz et al., 2015), we describe the orientations by a normal distribution function (NDF) $D(n)$ and its visible normal distribution function (VNDF) $D_{\omega_{o}}(n)$ when conditioned by a viewing direction $\omega_{o}$:
@@
The base surface BSDF can be arbitrary valid BSDF with standard evaluation and importance sampling procedures. In this work, we primarily work with the Disney principled BSDF (Burley, 2015) due to its capability to model a wide range of materials. The Disney principled BSDF mainly consists of a microfacet specular component and a diffuse component with empirical retroflection. We refer readers to Burley (2015) for the full definition. In Fig. 15, we show an example using an alternative thin-surface translucent base BSDF. We will also make the assumption that scattering in a primitive can be modeled accurately without an exitant positional distribution. See §9 for further discussion.
@@
We follow the original importance sampling strategy by (Heitz et al., 2015). Given a view direction $\omega_{o}$, we first sample the VNDF to generate a sample $n$. We then sample an incident direction $\omega_{i}$ using the existing strategy of the base BSDF. Since the VNDF sampling for SGGX is perfect, the sample weight is simply that of the base BSDF sampling. In practice, we spawn an indirect ray from the center of intersection segment and disable self-intersection. This is similar to common practice of sampling a Bidirectional Curve Scattering Distribution Function (BCSDF) in hair rendering.
@@ 7.2. Transmittance Optimization
[…]
@@ 8. Applications and Results
In Fig. 1, we demonstrate the versatility of our Gaussian primitive to represent objects with a wide range of geometric and material characteristics. The Dressing Table scene is modeled entirely by our primitives and contains parts that are acquired in different ways. The room, table, dragon, and logo are converted from meshes, while the plants, candle set, and blanket are converted from 3DGS reconstruction. Additionally, several objects, such as t he mirror and neon light bars, are modeled analytically. Our Gaussian primitives can adapt to different shapes, including flat surfaces and thin fibers, thanks to their anisotropic definition. The volumetric formulation naturally handles the fuzzy appearance from dense, stochastic details.
@@
Crucially, our framework supports full unbiased global illumination with Monte Carlo (volumetric) path tracing. This is in contrast to rasterization approaches including 3DGS, and volumetric ray marching approaches such as NeRF. Both families of approaches are limited to the first few dimensions of the path space and cannot solve the infinite dimensional light transport integral. Fig. 1 includes two renders with drastically different lighting setups. In the “daylight” setup, the scene is lit by an area light and an environment light; in the “night” setup, the scene is lit by a logo, neon light bars, and candles, all modeled as emissive Gaussian primitives. The number of emissive primitives exceeds 10K combined. Both renders showcase various global illumination effects including soft shadows, color bleeding, and inter-reflection. Additional sequences with zoomed-in camera animations are provided in the supplemental video.
@@
Comparison between Transmittance Models
In Fig. 12, we compare the rendering results using both our linear transmittance and the traditional exponential transmittance, where each Gaussian primitive defines a fraction of the extinction coefficient (Eq. 4) to be summed together. The exponential variants are obtained through a conversion process similar to that described in §7. They are rendered by brute-force analog decomposition tracking (Kutz et al., 2017) (Appendix D). We also show the differences between those renders and renders using the original scenes, while acknowledging that our method does not include a full inverse rendering pipeline to match the original renders exactly.
The selected scenes in Fig. 12 have different geometric characteristics. For the Checkerboards and the Office Chair that consist of flat surfaces (the chair is textured by a high-resolution opacity mask), the linear model clearly excels at producing opaque appearance with sharp silhouettes and shadows. The Plant and the Color Tree contain dense, unstructured elements. This type of scenes can be abstracted as exponential volumes because at far field, with enough sample size, the average free-flight distribution through these elements approaches exponential. However, this is not necessarily true when we represent the scenes as Gaussian primitives. To produce reasonable details, typically each primitive only covers one or a cluster of well aligned opaque elements, which is the case for those two scenes. At this granularity, the geometries inside a primitive still exhibit non-negligible spatial correlation. Therefore, the linear model is roughly on par with the exponential model for those two scenes. The exponential model could be more suitable for level-of-detail applications where a scene is represented as a low-resolution volume and viewed from afar, or true participating media content such as clouds and smoke. Finally, the quality improves with more primitives regardless of the underlying transmittance model.
Checkerboards Office Chair Plant Color Tree
Original
Refer to caption Refer to caption Refer to caption Refer to caption
Linear 5K PSNR: 28.31 Linear 50K PSNR: 30.05 Linear 10K PSNR: 26.80 Linear 20K PSNR: 21.28
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Exponential 5K PSNR: 25.10 Exponential 50K PSNR: 28.03 Exponential 10K PSNR: 26.59 Exponential 20K PSNR: 21.10
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Linear 25K PSNR: 29.88 Linear 250K PSNR: 30.44 Linear 50K PSNR: 26.86 Linear 100K PSNR: 22.95
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Exponential 25K PSNR: 26.43 Exponential 250K PSNR: 28.41 Exponential 50K PSNR: 27.44 Exponential 100K PSNR: 22.70
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 12. Comparison between linear and exponential transmittance model with fixed number of Gaussians. Error maps and PSNRs with respect to renders of the original scenes are provided. The linear model helps avoid leaking and produce sharp silhouettes, making it favorable in most cases we focus on. Even for scenes with dense elements such as the Plant and the Color Tree, each Gaussian usually only covers one or a few opaque elements, making the exponential model less suitable at this granularity.
@@
In Fig. 13, we compare our representation to the traditional volume representation consisting of regular voxels. We perform a simple voxelization of our models by resampling the free-flight PDF. For each voxel, we evaluate all overlapping Gaussians at the voxel center. The phase function parameters are similarly resampled. Only non-empty voxels are stored in a sparse voxel grid. The voxel grid is then linearly interpolated and rendered by ray marching. We evaluate the reconstruction quality using different voxel resolutions. Even using 8$\times$ more non-empty voxels than the number of Gaussians, the reconstruction quality is still significantly inferior. This is expected because unlike Gaussian mixtures that can approximate signals at arbitrary frequency, regular grid sampling is limited by the Nyquist-Shannon sampling theorem and the resolution must be at least twice the signal bandwidth to avoid aliasing. It would require an impractical amount of storage to properly represent the thin structures common in vegetation. Conversely, Gaussian primitives are much more effective at capturing fine geometry details.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 13. Comparison between Gaussian models and voxelized models at different resolutions. Only non-empty voxels are counted. Gaussian primitives are much more effective at representing sharp, detailed structure than voxels. Note that our current simple voxelization causes self-shadowing that leads to slight darkening. Direct illumination only.
@@
Garden Refer to caption Refer to caption Refer to caption
Stump Refer to caption Refer to caption Refer to caption
(a) Original (b) Lighting 1 (c) Lighting 2
@@
Refer to caption Refer to caption
Front view Back view
Figure 15. Our phase function is compatible with different base BSDFs. Here we show a leaf represented by our primitives and a translucent thin-surface base BSDF.